The temperature of 350 g of water changed from 20oC to 30 oC. How much heat did this sample
absorb? Cp for water = 4.18 J/goC
1
Expert's answer
2020-03-25T11:00:56-0400

The amount of heat needed to increasing of temperature of the body


"Q=cm\Delta T=4.18\times 350\times (30-20)=14630\:\rm J."